# EMOE: A Framework for Out-of-distribution Uncertainty Based Rejection via Model-Agnostic Expansive Matching of Experts

### v3

@@ 4.3 Ablation Studies
EMOE 96.64 96.15 94.15 93.81
Despite its dependency on base experts for pseudo-label generation, EMOE consistently improves performance across a wide range of base model types, including XGBoost, Random Forest, Decision Tree, and the OOD generalization method D-BAT (Pagliardini et al. 2023). As shown in Tab.4 (see details on this ablation in Appx.B.6). EMOE consistently improves upon each kind of base model, demonstrating its effectiveness as a model-agnostic framework. Moreover, this shows the future promise of incorporating EMOE with complementary methodology for OOD generalizability.
Per-head Matching Ablation Next, we ablate on the per-expert matching loss scheme (eq.7) used in EMOE with a simpler alternative: mean-only matching (MM) (eq.8) on the expanded points. This ablation tests whether the diversity induced by per-head supervision aids performance. Here we utilized single headed (SH) and multi-headed (MH) MLPs with mean matching, where EMOE uses multi-headed MLP with per expert loss (see  Mat.B.7 for details on the alternative losses). EMOE achieves the highest AUPRC improvements ($\Delta$) over base experts (+1.11 at R<<.2, +1.03 at R<<1) compared to SH+MM (+0.59, +0.45) and MH+MM (+0.25, +0.55), showing evidence for the effectiveness of training our embedding through the diverse multi-task loss induced by per-head matching ($\S$ 3).
Bottleneck Ablation  We motivated EMOE’s performance in terms of a multi-task scheme, which bottlenecks (the last hidden layer) to learn an embedding that can generalize to mimic predictions on a diverse set of pseudo-labels ($\S$ 3). We test this motivation by comparing the original architecture (Full: $2\times 512$ hidden layer) with a stronger bottleneck still, a small architecture (Tiny: $2\times 32$ hidden layer). (See Appx. B.8 for more details and results.) The performance gap is marginal between the ‘Full’ and ‘Tiny’ model (a $0.85\%$ difference) when using our proposed loss. In contrast, when using empirical risk minimization, we see a $4.87$ times bigger drop in performance between ‘Full’ and ‘Tiny’ models. This suggests that the bottlenecking properties of our method are key to EMOE’s performance, and show promise for EMOE in resource-constrained settings.
@@ 4.4 Discussion
Figure 6: Experts correlations between EMOE base experts and EMOE network heads on “Ames” dataset. (a) EMOE experts have a high correlation with EMOE base experts on sample the base experts make correct predictions. (b) EMOE experts show low correlations with EMOE base experts on samples where the base experts make incorrect predictions.
EMOE addresses the challenging single source OOD generalization problem for real-valued vector data, outperforming baselines, including those that leverage unlabelled OOD data. Across all metrics/datasets, EMOE achieves leading tally of 29, far exceeding both EMOE base and the next highest score of 4 from the strong competing baseline D-Bat (Pagliardini et al. 2023) as shown in Fig. 1. Despite its performance, EMOE does not incur an out-sized computational cost; the training of base experts may be done in parallel, and the cost of expanding data in the latent space is negligible (see Appx. A.3 for addition timing details).
In summary, EMOE presents a simple yet powerful model-agnostic approach to OOD generalization under single-source setting via pseudo-labels and expert-specific matching. EMOE’s architecture promotes a robust embedding and stands out in high confidence areas for OOD predictions. We believe this work opens promising directions for general-purpose, modality- and model- agnostic OOD learning, particularly in high-stakes applications like drug discovery, where confident extrapolation is critical and labeled data is scarce.

### v4

@@ 4.3 Ablation Studies
EXPLOR 96.64 96.15 94.15 93.81
Despite its dependency on base experts for pseudo-label generation, EXPLOR consistently improves performance across a wide range of base model types, including XGBoost, Random Forest, Decision Tree, and the OOD generalization method D-BAT (Pagliardini et al. 2023). As shown in Tab.4 (see details on this ablation in Appx.B.6). EXPLOR consistently improves upon each kind of base model, demonstrating its effectiveness as a model-agnostic framework. Moreover, this shows the future promise of incorporating EXPLOR with complementary methodology for OOD generalizability.
Per-head Matching Ablation Next, we ablate on the per-expert matching loss scheme (eq.7) used in EXPLOR with a simpler alternative: mean-only matching (MM) (eq.8) on the expanded points. This ablation tests whether the diversity induced by per-head supervision aids performance. Here we utilized single headed (SH) and multi-headed (MH) MLPs with mean matching, where EXPLOR uses multi-headed MLP with per expert loss (see  Mat.B.7 for details on the alternative losses). EXPLOR achieves the highest AUPRC improvements ($\Delta$) over base experts (+1.11 at R<<.2, +1.03 at R<<1) compared to SH+MM (+0.59, +0.45) and MH+MM (+0.25, +0.55), showing evidence for the effectiveness of training our embedding through the diverse multi-task loss induced by per-head matching ($\S$ 3).
Bottleneck Ablation  We motivated EXPLOR’s performance in terms of a multi-task scheme, which bottlenecks (the last hidden layer) to learn an embedding that can generalize to mimic predictions on a diverse set of pseudo-labels ($\S$ 3). We test this motivation by comparing the original architecture (Full: $2\times 512$ hidden layer) with a stronger bottleneck still, a small architecture (Tiny: $2\times 32$ hidden layer). (See Appx. B.8 for more details and results.) The performance gap is marginal between the ‘Full’ and ‘Tiny’ model (a $0.85\%$ difference) when using our proposed loss. In contrast, when using empirical risk minimization, we see a $4.87$ times bigger drop in performance between ‘Full’ and ‘Tiny’ models. This suggests that the bottlenecking properties of our method are key to EXPLOR’s performance, and show promise for EXPLOR in resource-constrained settings.
@@ 4.4 Discussion
Figure 6: Experts correlations between EXPLOR base experts and EXPLOR network heads on “Ames” dataset. (a) EXPLOR experts have a high correlation with EXPLOR base experts on sample the base experts make correct predictions. (b) EXPLOR experts show low correlations with EXPLOR base experts on samples where the base experts make incorrect predictions.
EXPLOR addresses the challenging single source OOD generalization problem for real-valued vector data, outperforming baselines, including those that leverage unlabelled OOD data. Across all metrics/datasets, EXPLOR achieves leading tally of 29, far exceeding both EXPLOR base and the next highest score of 4 from the strong competing baseline D-Bat (Pagliardini et al. 2023) as shown in Fig. 1. Despite its performance, EXPLOR does not incur an out-sized computational cost; the training of base experts may be done in parallel, and the cost of expanding data in the latent space is negligible (see Appx. A.3 for addition timing details).
In summary, EXPLOR presents a simple yet powerful model-agnostic approach to OOD generalization under single-source setting via pseudo-labels and expert-specific matching. EXPLOR’s architecture promotes a robust embedding and stands out in high confidence areas for OOD predictions. We believe this work opens promising directions for general-purpose, modality- and model- agnostic OOD learning, particularly in high-stakes applications like drug discovery, where confident extrapolation is critical and labeled data is scarce.